# Distributionally and Adversarially Robust Logistic Regression via Intersecting Wasserstein Balls

### v2

@@ 1 INTRODUCTION
Supervised learning traditionally involves access to a training dataset whose instances are assumed to be independently sampled from a true data-generating distribution (Bishop,, 2006; Hastie et al.,, 2009). Optimizing an expected loss for the empirical distribution constructed from such a training set, also known as empirical risk minimization (ERM), enjoys several desirable properties in relatively generic settings, including convergence to the true risk minimization problem as the number of training samples increases (Vapnik,, 1999, Chapter 2). In real-world applications, however, various challenges, such as data scarcity and the existence of adversarial attacks, lead to deteriorated out-of-sample performance for models trained via ERM.
@@
One of the key limitations of ERM, particularly as it is designed to minimize an expected loss for the empirical distribution, emerges from the finite nature of data in practice. This leads ERM to suffer from the ‘optimism bias’, also known as overfitting (Murphy,, 2022), or the optimizer’s curse (DeMiguel and Nogales,, 2009; Smith and Winkler,, 2006), causing deteriorated out-of-sample performance. A popular approach to prevent this phenomenon, distributionally robust optimization (DRO; Delage and Ye, 2010), optimizes the expected loss for the worst-case distribution residing within a pre-specified ambiguity set.
@@
Another key challenge faced by ERM in practice is adversarial attacks, where an adversary perturbs the observed features during the testing or deployment phase (Szegedy et al.,, 2014; Goodfellow et al.,, 2015). For neural networks, the paradigm of adversarial training (AT; Madry et al., 2018) is thus designed to provide adversarial robustness by simulating the attacks in the training stage. Several successful variants of AT, specialized to different losses and attacks, have been proposed in the literature to achieve adversarial robustness without significantly reducing performance on training sets (Shafahi et al.,, 2019; Zhang et al.,, 2019; Pang et al.,, 2022; Gao et al.,, 2019). While some works (e.g., Carlini et al., 2019; Uesato et al., 2018) examine adversarial robustness guarantees of various training algorithms, a recent stream of research (e.g., Bennouna et al., 2023; Xing et al., 2022b ) focuses on adversarially robust optimization (ARO), constraining ERM to guarantee an exact pre-specified level of adversarial robustness while maximizing training accuracy.
@@
Recently, it has been observed that the two aforementioned notions of robustness can conflict, as adversarially robust (AR) models suffer from severe overfitting (robust overfitting; Raghunathan et al., 2019; Yu et al., 2022; Li and Spratling, 2023). Indeed, it is observed that robust overfitting is even more severe than traditional overfitting (Rice et al.,, 2020). To this end, some works address robust overfitting by revisiting AT algorithms and adding adjustments for better generalization (Chen et al.,, 2020; Li and Li,, 2023). In a recent work, Bennouna et al., (2023, Thm 3.2) decompose the error gap of robust overfitting into the statistical error of estimating the true data-generating distribution via the empirical distribution and an adversarial error resulting from the adversarial attacks, hence proposing the simultaneous adoption of DRO and ARO.
@@
In this work, we study logistic regression (LR) for binary classification that is adversarially robust against $\ell_{p}$-attacks (Croce et al.,, 2020). To address robust overfitting faced by the adversarially robust LR model, we employ a DRO approach where distributional ambiguity is modeled with the type-1 Wasserstein metric. We base our work on an observation that the worst-case logistic loss under adversarial attacks can be represented as a Lipschitz continuous and convex loss function. This allows us to use existing Wasserstein DRO machinery for Lipschitz losses, and derive an exact reformulation of the Wasserstein DR counterpart of adversarially robust LR as a tractable convex problem.
@@
Our main contribution lies in reducing the size of the Wasserstein ambiguity set in the DRO problem mentioned above, in order to create a less conservative problem while preserving the same robustness guarantees. To accomplish this, we draw inspiration from recent work on ARO that leverages auxiliary datasets (e.g., Gowal et al., 2021; Xing et al., 2022b ) and revise our DRO problem by intersecting its ambiguity set with another ambiguity set constructed using an auxiliary dataset. Examples of auxiliary data include synthetic data generated from a generative model (e.g., privacy-preserving data release), data in the presence of distributional shifts (e.g., different time periods/regions), noisy data (e.g., measurement errors), or out-of-domain data (e.g., different source); any auxiliary dataset is viable as long as its instances are sampled independently from an underlying data-generating distribution whose Wasserstein distance to the true data-generating distribution is known or can be estimated. Figure 1 illustrates our framework.
@@ 2 RELATED WORK
Auxiliary data in ARO. The use of auxiliary data appears in the ARO literature. In particular, it is shown that additional unlabeled data sampled from the same (Carmon et al.,, 2019; Xing et al., 2022a, ) or different (Deng et al.,, 2021) data-generating distributions could provide adversarial robustness. Sehwag et al., (2022) show that adversarial robustness can be certified even when it is provided for a synthetic dataset as long as the distance between its generator and the true data-generating distribution can be quantified. Gowal et al., (2021); Xing et al., 2022b propose optimizing a weighted combination of ARO over empirical and synthetic datasets. We show that the latter approach can be recovered by our model.
@@
DRO-ARO interactions. In our work, we optimize ARO against worst-case data-generating distributions residing in an ambiguity set, where the type-1 Wasserstein metric is used for distances since it is arguably the most common choice in machine learning (ML) with Lipschitz losses (Shafieezadeh-Abadeh et al.,, 2019; Gao,, 2023). In the literature, it is shown that standard ARO is equivalent to the DRO of the original loss function with a type-$\infty$ Wasserstein metric (Staib and Jegelka,, 2017; Khim and Loh,, 2018; Pydi and Jog,, 2021; Regniez et al.,, 2022; Frank and Niles-Weed,, 2024), or a Lévy-Prokhorov metric (Bennouna and Van Parys,, 2022). In other words, in the absence of adversarial attacks, training models adversarially with artificial attacks provide some distributional robustness. Hence, our DR ARO approach can be interpreted as optimizing the logistic loss over the worst-case distribution whose $1$-Wasserstein distance is bounded by a pre-specified radius from at least one distribution residing in an $\infty$-Wasserstein ball around the empirical distribution. Conversely, Sinha et al., (2018) discuss that while DRO over Wasserstein balls is intractable for generic losses (e.g., neural networks), its Lagrange relaxation resembles ARO and thus ARO yields a certain degree of (relaxed) distributional robustness (Wu et al.,, 2020; Bui et al.,, 2022; Phan et al.,, 2023). However, to the best of our knowledge, there have not been works optimizing a pre-specified level of type-$1$ Wasserstein distributional robustness (that hedges against overfitting, Kuhn et al., 2019) and adversarial robustness (that hedges against adversarial attacks, Goodfellow et al., 2015) simultaneously. To our knowledge, the only approach that considers the exact DR counterpart of ARO is proposed by Bennouna et al., (2023) who model distributional ambiguity with $\varphi$-divergences for neural networks.
@@
Intersecting ambiguity sets in DRO. Recent work started to explore the intersection of ambiguity sets for different contexts (Awasthi et al.,, 2022; Wang et al.,, 2024) or different metrics (Zhang et al.,, 2023). Our idea of intersecting Wasserstein balls is originated from the “Surround, then Intersect” strategy (Taskesen et al.,, 2021, §5.2) to train linear regression under sequential domain adaptation in a non-adversarial setting (see Shafahi et al., 2020 and Song et al., 2019 for robustness in domain adaptation/transfer learning). The aforementioned work focuses on the squared loss function with an ambiguity set using the Wasserstein metric developed for the first and second distributional moments. In a recent study, Rychener et al., (2024) generalize most of the previous results and prove that DRO problems over the intersection of two Wasserstein balls admit tractable convex reformulations whenever the loss function is the maximum of concave functions
@@
Logistic loss in DRO and ARO. Our choice of LR aligns with the current directions and open questions in the related literature. In the DRO literature, even in the absence of adversarial attacks, the aforementioned work of Taskesen et al., (2021) on the intersection of Wasserstein ambiguity sets is restricted to linear regression. The authors show that this problem admits a tractable convex optimization reformulation, and their proof relies on the properties of the squared loss. Similarly, Rychener et al., (2024) discuss that the logistic loss fails to satisfy the piece-wise concavity assumption and is inherently difficult to optimize over the intersection of Wasserstein balls. We contribute to the DRO literature for adversarial and non-adversarial settings because we show that such a problem would be NP-hard for the logistic loss even without adversarial attacks, and develop specialized approximation techniques. Our problem recovers DR LR (Shafieezadeh-Abadeh et al.,, 2015; Selvi et al.,, 2022) as a special case in the absence of adversarial attacks and auxiliary data. Answering theoretical challenges posed by logistic regression has been useful in answering more general questions in the DRO literature, such as DR LR (Shafieezadeh-Abadeh et al.,, 2015) leading to DR ML (Shafieezadeh-Abadeh et al.,, 2019) and mixed-feature DR LR (Selvi et al.,, 2022) leading to mixed-feature DR Lipschitz ML (Belbasi et al.,, 2023). Finally, in the (non-DR) ARO literature, there are recent theory developments on understanding the effect of auxiliary data (e.g.Xing et al., 2022b ) specifically for squared and logistic loss functions.
@@ 3 PRELIMINARIES
The distance between instances $\bm{\xi}=(\bm{x},y)\in\Xi$ and $\bm{\xi^{\prime}}=(\bm{x^{\prime}},y^{\prime})\in\Xi$ for $\kappa>0$ and $q\geq 1$ is
$d(\bm{\xi},\bm{\xi^{\prime}})=\lVert\bm{x}-\bm{x^{\prime}}\rVert_{q}+\kappa\cdot\mathds{1}{[y\neq y^{\prime}]}.$
@@
where $\mathfrak{B}_{\varepsilon}(\mathbb{P}):=\{\mathbb{Q}\in\mathcal{P}(\Xi):\mathrm{W}(\mathbb{Q},\mathbb{P})\leq\varepsilon\}$ denotes the Wasserstein ball centered at $\mathbb{P}\in\mathcal{P}(\Xi)$ with radius $\varepsilon$. We refer to Mohajerin Esfahani and Kuhn, (2018) and Kuhn et al., (2019) for the properties of 6.
@@
The goal of adversarial robustness is to provide robustness against adversarial attacks (Goodfellow et al.,, 2015). An adversarial attack, in the widely studied $\ell_{p}$-noise setting (Croce et al.,, 2020), perturbs the features of the test instances $(\bm{x},y)$ by adding additive noise $\bm{z}$ to $\bm{x}$. The adversary chooses the noise vector $\bm{z}$, subject to $\lVert\bm{z}\rVert_{p}\leq\alpha$, so as to maximize the loss $\ell_{\bm{\beta}}(\bm{x}+\bm{z},y)$ associated with this perturbed test instance. Therefore, ARO solves the following optimization problem in the training stage to hedge against adversarial perturbations at the test stage:
@@
8 reduces to 4 when $\alpha=0$. Note that 8 is identical to feature robust training (Bertsimas et al.,, 2019) which is not motivated by adversarial attacks, but the presence of noisy observations in the training set (Ben-Tal et al.,, 2009; Gorissen et al.,, 2015).
DRO-ARO connection. A connection between ARO and DRO is noted in the literature (Staib and Jegelka, 2017, Proposition 3.1, Khim and Loh, 2018, Lemma 22, Pydi and Jog, 2021, Lemma 5.1, Regniez et al., 2022, Proposition 2.1, Frank and Niles-Weed, 2024, Lemma 3, and Bennouna et al., 2023, §3). Namely, problem 8 is equivalent to a DRO problem
@@
where the ambiguity set $\mathfrak{B}^{\infty}_{\alpha}({\mathbb{P}}_{N})$ is a type-$\infty$ Wasserstein ball (Givens and Shortt,, 1984) with radius $\alpha$. Hence, in non-adversarial settings, 8 provides robustness with respect to the type-$\infty$ Wasserstein distance. In the case of adversarial attacks, it suffers from robust overfitting as discussed earlier. To address this issue, one straightforward approach is to revisit (10) and replace $\alpha$ with some $\alpha^{\prime}>\alpha$. This approach, however, does not provide improvements for the out-of-sample performance since (i) the type-$\infty$ Wasserstein distance employed in problem (10) uses a metric on the feature space, ignoring labels; (ii) type-$\infty$ Wasserstein distances do not provide strong out-of-sample performances in ML (unlike, e.g., the type-$1$ Wasserstein distance) since the required radii to provide meaningful robustness guarantees are typically too large (Bennouna and Van Parys,, 2022, §1.2.2, and references therein). We thus study the type-$1$ Wasserstein counterpart of 8, which we initiate in the next section.

### v4

@@ 1 INTRODUCTION
Supervised learning traditionally involves access to a training dataset whose instances are assumed to be independently sampled from a true data-generating distribution (Bishop, 2006; Hastie et al., 2017). Optimizing an expected loss for the empirical distribution constructed from such a training set, also known as empirical risk minimization (ERM), enjoys several desirable properties in relatively generic settings, including convergence to the true risk minimization problem as the number of training samples increases (Vapnik, 1999, Chapter 2). In real-world applications, however, various challenges, such as data scarcity and the existence of adversarial attacks, lead to deteriorated out-of-sample performance for models trained via ERM.
@@
One of the key limitations of ERM, particularly as it is designed to minimize an expected loss for the empirical distribution, emerges from the finite nature of data in practice. This leads ERM to suffer from the ‘optimism bias’, also known as overfitting (Murphy, 2022), or the optimizer’s curse (DeMiguel and Nogales, 2009; Smith and Winkler, 2006), causing deteriorated out-of-sample performance. A popular approach to prevent this phenomenon, distributionally robust optimization (DRO; Delage and Ye 2010), optimizes the expected loss for the worst-case distribution residing within a pre-specified ambiguity set.
@@
Another key challenge faced by ERM in practice is adversarial attacks, where an adversary perturbs the observed features during the testing or deployment phase (Szegedy et al., 2014; Goodfellow et al., 2015), also known as evasion corruption at test time (Biggio et al., 2013). For neural networks, the paradigm of adversarial training (AT; Madry et al. 2018) is thus designed to provide adversarial robustness by simulating such attacks in the training stage. Several successful variants of AT, specialized to different losses and attacks, have been proposed in the literature to achieve adversarial robustness without significantly reducing performance on training sets (Shafahi et al., 2019; Zhang et al., 2019; Gao et al., 2019; Pang et al., 2022). Some studies (Uesato et al., 2018; Carlini et al., 2019; Wu et al., 2020) investigate the adversarial robustness guarantees of various training algorithms, leading to a research direction focused on heuristic improvements to such models (e.g., Rade and Moosavi-Dezfooli 2022). Our work aligns with another recent line of research (Xing et al., 2022a; Bennouna et al., 2023) on adversarially robust optimization (ARO), which constrains ERM to guarantee an exact, pre-specified level of adversarial robustness while maximizing training accuracy.
@@
Recently, it has been observed that the two aforementioned notions of robustness can be at odds, as adversarially robust (AR) models suffer from severe overfitting (robust overfitting; Raghunathan et al. 2019; Yu et al. 2022; Li and Spratling 2023). Indeed, it is observed that robust overfitting is even more severe than traditional overfitting (Rice et al., 2020). To this end, some works address robust overfitting by revisiting AT algorithms and adding adjustments for better generalization (Chen et al., 2020; Li and Li, 2023). In a recent work, Bennouna et al. (2023, Thm 3.2) decompose the error gap of robust overfitting into the statistical error of estimating the true data-generating distribution via the empirical distribution and an adversarial error resulting from the adversarial attacks, hence proposing the simultaneous adoption of DRO and ARO.
@@
In this work, we study logistic regression (LR) for binary classification that is adversarially robust against $\ell_{p}$-attacks (Croce et al., 2020). To address robust overfitting faced by the adversarially robust LR model, we employ a DRO approach where distributional ambiguity is modeled with the type-1 Wasserstein metric. We base our work on an observation that the worst-case logistic loss under adversarial attacks can be represented as a Lipschitz continuous and convex loss function. This allows us to use existing Wasserstein DRO machinery for Lipschitz losses, and derive an exact reformulation of the Wasserstein DR counterpart of adversarially robust LR as a tractable convex problem.
@@
Our main contribution lies in reducing the size of the Wasserstein ambiguity set in the DRO problem mentioned above, in order to create a less conservative problem while preserving the same distributional robustness guarantees. To accomplish this, we draw inspiration from recent work on ARO that leverages auxiliary datasets (e.g., Gowal et al. 2021; Xing et al. 2022a) and revise our DRO problem by intersecting its ambiguity set with another ambiguity set constructed using an auxiliary dataset. Examples of auxiliary data include synthetic data generated from a generative model (e.g., privacy-preserving data release), data in the presence of distributional shifts (e.g., different time periods/regions), noisy data (e.g., measurement errors), or out-of-domain data (e.g., different source); any auxiliary dataset is viable as long as its instances are sampled independently from an underlying data-generating distribution whose Wasserstein distance to the true data-generating distribution is known or can be estimated. Figure 1 illustrates our framework.
@@ 2 RELATED WORK
Auxiliary data in ARO. The use of auxiliary data appears in the ARO literature. In particular, it is shown that additional unlabeled data sampled from the same (Carmon et al., 2019; Xing et al., 2022b) or different (Deng et al., 2021) data-generating distributions could provide adversarial robustness. Sehwag et al. (2022) show that adversarial robustness can be certified even when it is provided for a synthetic dataset as long as the distance between its generator and the true data-generating distribution can be quantified. Gowal et al. (2021) and Xing et al. (2022a) propose optimizing a weighted combination of ARO over empirical and synthetic datasets. We show that the latter approach can be recovered by our model.
@@
DRO-ARO interactions. In this work, we optimize ARO against worst-case data-generating distributions residing in an ambiguity set, where the type-1 Wasserstein metric is used for distances since it is arguably the most common choice in machine learning (ML) with Lipschitz losses (Shafieezadeh-Abadeh et al., 2019; Gao, 2023). In the literature, it is shown that standard ARO is equivalent to the DRO of the original loss function with a type-$\infty$ Wasserstein metric (Staib and Jegelka, 2017; Khim and Loh, 2018; Pydi and Jog, 2021; Regniez et al., 2022; Frank and Niles-Weed, 2024). In other words, in the absence of adversarial attacks, training models adversarially with artificial attacks provide some distributional robustness. Hence, our DR ARO approach can be interpreted as optimizing the logistic loss over the worst-case distribution whose $1$-Wasserstein distance is bounded by a pre-specified radius from at least one distribution residing in an $\infty$-Wasserstein ball around the empirical distribution. Conversely, Sinha et al. (2018) discuss that while DRO over Wasserstein balls is intractable for generic losses (e.g., neural networks), its Lagrange relaxation resembles ARO and thus ARO yields a certain degree of (relaxed) distributional robustness (Wu et al., 2020; Bui et al., 2022; Phan et al., 2023). Such literature suggests that, when there is no concern about the statistical errors caused by using empirical distributions (e.g., in very high-data regimes), one can train DR models to obtain adversarial robustness guarantees. However, as discussed by Bennouna and Van Parys (2022), when statistical errors exist, then we need to be simultaneously robust against adversarial attacks and statistical errors. To the best of our knowledge, there have not been works optimizing a pre-specified level of type-$1$ Wasserstein distributional robustness (that hedges against overfitting, Kuhn et al. 2019) and adversarial robustness (that hedges against adversarial attacks, Goodfellow et al. 2015) simultaneously. To our knowledge, the only approach that considers the exact DR counterpart of ARO is proposed by Bennouna et al. (2023), who model distributional ambiguity with $\varphi$-divergences for neural networks.
@@
Intersecting ambiguity sets in DRO. Recent work started to explore the intersection of ambiguity sets for different contexts (Awasthi et al., 2022; Wang et al., 2024) or different metrics (Tanoumand et al., 2023). Our idea of intersecting Wasserstein balls is originated from the “Surround, then Intersect” strategy (Taskesen et al., 2021, §5.2) to train linear regression under sequential domain adaptation in a non-adversarial setting (see the work of Shafahi et al. 2020 and Song et al. 2019 for robustness in domain adaptation/transfer learning). The aforementioned work focuses on the squared loss function with an ambiguity set using the Wasserstein metric developed for the first and second distributional moments. In a recent study, Rychener et al. (2024) generalize most of the previous results and prove that DRO problems over the intersection of two Wasserstein balls admit tractable convex reformulations whenever the loss function is the maximum of concave functions. They also discuss why distributions lying in the intersection of two Wasserstein balls are more natural candidates for the unknown true distribution than those that are Wasserstein barycenters or mixture distributions of the empirical and auxiliary distributions (referred to as heterogeneous data sources; see Example 1, Proposition 2, and Corollary 1).
@@
Logistic loss in DRO and ARO. Our choice of LR aligns with the current directions and open questions in the related literature. In the DRO literature, even in the absence of adversarial attacks, the aforementioned work of Taskesen et al. (2021) on the intersection of Wasserstein ambiguity sets is restricted to linear regression. The authors show that this problem admits a tractable convex optimization reformulation, and their proof relies on the properties of the squared loss. Similarly, Rychener et al. (2024) discuss that the logistic loss fails to satisfy the piece-wise concavity assumption and is inherently difficult to optimize over the intersection of Wasserstein balls. We contribute to the DRO literature for adversarial and non-adversarial settings because we show that such a problem would be NP-hard for the logistic loss even without adversarial attacks, and develop specialized approximation techniques. Our problem recovers DR LR (Shafieezadeh-Abadeh et al., 2015; Selvi et al., 2022) as a special case in the absence of adversarial attacks and auxiliary data. Answering theoretical challenges posed by logistic regression has been useful in answering more general questions in the DRO literature, such as DR LR (Shafieezadeh-Abadeh et al., 2015) leading to DR ML (Shafieezadeh-Abadeh et al., 2019) and mixed-feature DR LR (Selvi et al., 2022) leading to mixed-feature DR Lipschitz ML (Belbasi et al., 2023). Finally, in the (non-DR) ARO literature, there are recent theory developments on understanding the effect of auxiliary data (e.g.Xing et al. 2022a) specifically for squared and logistic loss functions.
Single-step adversarial training and single-shot ARO.
Our work proposes a single-shot convex optimization procedure to train logistic models that are both adversarially and distributionally robust. Although the terminology may resemble the recent work on single-step adversarial training for neural networks (Wong et al., 2020; Lin et al., 2023, 2024), the two approaches operate differently. Single-step adversarial training generates adversarial perturbations using a single gradient computation at each iteration of iterative model training and improves robustness through updates, with performance typically assessed at intermediate checkpoints. In contrast, our method solves a convex optimization problem once to obtain a model that satisfies both forms of robustness by design. To enable tractability, this approach leverages the convexity and Lipschitz continuity of the loss function under adversarial attacks, which hold for the logistic loss function. While single-step adversarial training applies broadly to general classes of models such as neural networks, our framework offers a complementary, optimization-based perspective in the logistic regression model, where structural properties can be fully exploited.
@@ 3 PRELIMINARIES
The distance between instances $\bm{\xi}=(\bm{x},y)\in\Xi$ and $\bm{\xi^{\prime}}=(\bm{x^{\prime}},y^{\prime})\in\Xi$ for $\kappa\geq 0$ and $q\geq 1$ is
$d(\bm{\xi},\bm{\xi^{\prime}})=\lVert\bm{x}-\bm{x^{\prime}}\rVert_{q}+\kappa\cdot\mathds{1}\small{[y\neq y^{\prime}]}.$
@@
where $\mathfrak{B}_{\varepsilon}(\mathbb{P}):=\{\mathbb{Q}\in\mathcal{P}(\Xi):\mathrm{W}(\mathbb{Q},\mathbb{P})\leq\varepsilon\}$ denotes the Wasserstein ball centered at $\mathbb{P}\in\mathcal{P}(\Xi)$ with radius $\varepsilon$. We refer to Mohajerin Esfahani and Kuhn (2018) and Kuhn et al. (2019) for the properties of 6 and estimating $\varepsilon$.
@@
The goal of adversarial robustness is to provide robustness against adversarial attacks (Goodfellow et al., 2015). An adversarial attack, in the widely studied $\ell_{p}$-noise setting (Croce et al., 2020), perturbs the features of the test instances $(\bm{x},y)$ by adding additive noise $\bm{z}$ to $\bm{x}$. The adversary chooses the noise vector $\bm{z}$, subject to $\lVert\bm{z}\rVert_{p}\leq\alpha$, so as to maximize the loss $\ell_{\bm{\beta}}(\bm{x}+\bm{z},y)$ associated with this perturbed test instance. Therefore, ARO solves the following optimization problem in the training stage to hedge against adversarial perturbations at the test stage:
@@
8 reduces to 4 when $\alpha=0$. Note that 8 is identical to feature robust training (Bertsimas et al., 2019) which is not motivated by adversarial attacks, but by the presence of noisy observations in the training set (Ben-Tal et al., 2009; Gorissen et al., 2015).
DRO-ARO connection. A connection between ARO and DRO is noted in the literature (Staib and Jegelka 2017, Proposition 3.1, Khim and Loh 2018, Lemma 22, Pydi and Jog 2021, Lemma 5.1, Regniez et al. 2022, Proposition 2.1, Frank and Niles-Weed 2024, Lemma 3, and Bennouna et al. 2023, §3). Namely, problem 8 is equivalent to a DRO problem
@@
where the ambiguity set $\mathfrak{B}^{\infty}_{\alpha}({\mathbb{P}}_{N})$ is a type-$\infty$ Wasserstein ball (Givens and Shortt, 1984) with radius $\alpha$. Hence, in non-adversarial settings, 8 provides robustness with respect to the type-$\infty$ Wasserstein distance. In the case of adversarial attacks, it suffers from robust overfitting as discussed earlier. To address this issue, one straightforward approach is to revisit (10) and replace $\alpha$ with some $\alpha^{\prime}>\alpha$. This approach, however, does not provide improvements for the out-of-sample performance since (i) the type-$\infty$ Wasserstein distance employed in problem (10) uses a metric on the feature space, ignoring labels; (ii) type-$\infty$ Wasserstein distances do not provide strong out-of-sample performances in ML (unlike, e.g., the type-$1$ Wasserstein distance) since the required radii to provide meaningful robustness guarantees are typically too large (Bennouna and Van Parys, 2022, §1.2.2, and references therein). We thus study the type-$1$ Wasserstein counterpart of 8, which we initiate in the next section.